Find the equation of the locus of the center of a moving circle tangent to the y-axis and to a circle with a radius 2 with center at (8, 2).


1
Expert's answer
2021-10-18T15:56:03-0400

Let "P(x,y)" be the center of a moving circle tangent to the y-axis and to a circle with a radius 2 with center at (8, 2).

The distance from point "P" to the y-axis is "|x|". The distance from the point "P" to a center "(8, 2)" is

"\sqrt{(x-8)^2+(y-2)^2}." The distance from the point "P" to a circle is "\sqrt{(x-8)^2+(y-2)^2}-2."

Then

"\sqrt{(x-8)^2+(y-2)^2}-2=|x|"


"\sqrt{(x-8)^2+(y-2)^2}=|x|+2"

"x^2-16x+64+(y-2)^2=x^2+4|x|+4"

"(y-2)^2=16x+4|x|-60"

Since "(y-2)^2\geq0, y\in \R," we take "|x|=x, x\geq0"


"(y-2)^2=16x+4x-60"

"(y-2)^2=20x-60"

The equation of the locus of the center of a moving circle tangent to the y-axis and to a circle with a radius 2 with center at (8, 2) is the equaion of the parabola


"(y-2)^2=20x-60"